OpenStudy (chacha1234):

How many solutions does this system have? x-7=-4 3x+y=8

OpenStudy (elson):

x-7=-4 so x=3. Sub in 3 to the 3x+y=8. the solution is (3,-1) there is only one

OpenStudy (chacha1234):

How many solutions does this system have? 12x-4y=-8 y=3x+2

OpenStudy (elson):

Sub in y=3x+2. 12x-12y-8=-8. The 2 lines are the same

OpenStudy (elson):

Did that help?

OpenStudy (chacha1234):

So.. no solution...?

OpenStudy (elson):

Everything is a solution. If a point is on one of the lines, it is on the other line

OpenStudy (elson):

infinity

OpenStudy (chacha1234):

How many solutions does this system have? y=3x-5 y=-x+4

OpenStudy (elson):

got to go

OpenStudy (chacha1234):

One more please

OpenStudy (chacha1234):

Just that one